QUESTION 3 (25 marks) A bioengineer is evaluating methods used to apply adhesives onto microporous paper tape that is commonly use after surgery. The machinery costs RM800,000, has no salvage value, and the BTCF estimate is RM300,000 per year for up to 10 years. The effective income tax is 38% and after taxes MARR is 8% per year. The two depreciation methods to consider are: MACRS with n 5 years and Straight-line with n 8 years (neglect the half year convention effect). For a study period of 8 years: Which depreciation method offers the better tax advantage? Please explain and describe your answer briefly. a. (15 marks) b. Demonstrate that the same total taxes are paid for both depreciation methods (10 marks)
Answer #1

Answer:

Part A)

For MACRS depreciation

PV = ATCF/(1+MARR)^t

MARR=8%

t= year in which cash flow occur

For year 1

PV=24680/(1+8%)^1=$228518.52

MACRS Method for depreciation
Year 0 1 2 3 4 5 6 7 8
Initial Investment 800000
BTCF 300000 300000 300000 300000 300000 300000 300000 300000
Depreciation rate 20% 32% 19.20% 11.52% 11.52% 5.76% 0% 0%
Depreciation D 160000 256000 153600 92160 92160 46080 0 0
Taxable Income I= BTCF-D 140000 44000 146400 207840 207840 253920 300000 300000
Tax  @38% T=I*38% 53200 16720 55632 78979.2 78979.2 96489.6 114000 114000
ATCF =BTCF-T 246800 283280 244368 221020.8 221020.8 203510.4 186000 186000
PV of ATCF @ MARR=8% 228518.52 242866.94 193987.20 162456.89 150423.04 128246.07 108529.21 100490.01
NPV= Sum of PV of ATCF-Initial Investment 515517.88

For Straight line method

Depreciation rate=1/project life=12.5%

Straight Line Method for depreciation
Year 0 1 2 3 4 5 6 7 8
Initial Investment 800000
BTCF 300000 300000 300000 300000 300000 300000 300000 300000
Depreciation rate 12.50% 12.50% 12.50% 12.50% 12.50% 12.50% 12.50% 12.50%
Depreciation D 100000 100000 100000 100000 100000 100000 100000 100000
Taxbable Income I= BTCF-D 200000 200000 200000 200000 200000 200000 200000 200000
Tax  @38% T=I*38% 76000 76000 76000 76000 76000 76000 76000 76000
ATCF 224000 224000 224000 224000 224000 224000 224000 224000
PV of ATCF @ MARR=8% 207407.41 192043.90 177818.42 164646.69 152450.64 141158.00 130701.85 121020.23
NPV= Sum of PV of ATCF-Initial Investment 487247.12

Since NPV for MACRS depreciation method is greater than NPV for Straight line method.So MACRS method provides better tax advantage.

Part B)

Total Tax paid in MACRS method=53200+16720+55632+78979.2+78979.2+96489.6+114000+114000=$608000 eq 1

Total Tax paid in Straight Line method =76000+76000+76000+76000+76000+76000+76000+76000=$608000 eq 2

From above equation 1 and 2 it is clear that total tax paid in MACRS and Straight line is same.